Solve the following system by graphing.
x - y = 4
x + y = 2

Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

Solve the following system by graphing.
x - y = 4
x + y = 2
-----
Add and solve for "x"::
2x = 6
x = 3
----
Solve for "y"::
3 + y = 2
y = -1
-------
Cheers,
Stan H.
----------

Answer by MathTherapy(10552)


Solve the following system by graphing.
x - y = 4
x + y = 2
Set the x-value for the 1st equation, equal to 0. Solve for y to determine the y-intercept-point
Set the y-value for the 1st equation, equal to 0. Solve for x to determine the x-intercept-point
Join the two points
Repeat the process for the 2nd equation
The point of intersection for both lines will be the solution to the system.